I would appreciate your help in solving this nonlinear equation:
Problem: (x+2)^2+(y-2)^2=13 (1)
( 2x+y=6 (2)

Thank you greatly for all your help.

Answer by scott8148(6628)

use the second equation to find y (y=6-2x) and then substitute this value into the first equation

(x+2)^2+(6-2x-2)^2=13 ___ x^2+4x+4+16-16x+4x^2=13 ___ 5x^2-12x+7=0 ___ (5x-7)(x-1)=0

plug the x values back into the second equation to find the corresponding y values ___ there will be two pairs

this is a line intersecting a circle (at two points)